# Emergent predictability in microbial ecosystems

### v1

@@ Abstract
A long-standing hope of theoretical ecology is that, despite this complexity, some patterns might nevertheless be predictable. The search for emergent regularities in complex ecosystems is as old as ecology itself. For example, classical work uncovered reproducible trophic structures across taxonomically divergent communities [1] or a consistent ratio of predator species to prey species [2, 3]. More recently, moving beyond the patterns persisting despite high ecosystem diversity, it was proposed that some regularities could perhaps arise because of it, coining the term “emergent simplicity” [4, 5, 6, 7, 8]. This notion is at the foundation of the interface between statistical physics and the so-called large-$N$ ecology, explaining species abundance distributions and other patterns as some parameter becomes large [9, 10, 11, 12, 13, 14, 15, 16].
@@
Figure 1: A quantitative framework for interrogating emergent simplicity in data. A: A general experimental protocol with community richness (number of strains) as a control parameter. A fixed library of strains is subsampled to assemble communities of low richness (top) or high richness (bottom). Community composition (strain abundances, $n_{i}$) and some property of interest $Y$ are measured in each community. B: Our approach will be to train models predicting $Y$ based on combined abundances of groups of taxa. Each choice of a coarsening $\Psi$ defines a regression model (coefficients $\{\beta_{i}^{\Psi}\}$), whose performance we will evaluate. C: For each coarsening $\Psi$, we can evaluate the description complexity $I(\Psi)$ (the extent of coarsening; see text) and the out-of-sample prediction error, quantified as the fraction of variance unexplained (FVU) by the corresponding regression model. The cartoon illustrates the expected general shape of the scatter plot in these axes, evaluated for all possible coarsenings $\Psi$ (points). Our focus is on the shape of the Pareto front (thick line), characterizing the best predictive power achievable by models of a given complexity. As community richness increases, the phenomenon of emergent simplicity is identified by the behavior of the Pareto fronts, and can take one of two distinct forms: D: emergent coarse-grainability, or E: emergent predictability (vertical arrow). We stress that our normalization—focusing on the fraction of variance unexplained, equivalent to normalizing all observables to unit variance—ensures that both forms of emergent simplicity defined here are distinct from the simplest law-of-large-numbers intuition (a.k.a. central limit theorem or “self-averaging”); see text.
@@
The idea of emergent simplicity is very appealing, suggesting that functional predictions could perhaps be made directly in the complex natural context [17]. However, so far this hope remains out of reach. Most reported instances of “emergent simplicity” fall under the umbrella of functional convergence or self-averaging, focusing on reproducibility rather than prediction [1, 2, 3, 4, 5, 6, 7, 8]. Unfortunately, such reproducibility is often approximate or limited to simplified conditions [18], and even when it holds, offers limited predictive power. As an illustration, consider the statement: “if the soil ecosystem is diverse enough, the rate of process $P$ will be $\approx\!\!X$ per gram of biomass”. This blanket expectation would likely be somewhat naïve, but even if it were true, it would leave key questions unanswered – such as how the process rate would be affected by a perturbation, or what intervention would help increase it. As a result, observing emergent regularities of this kind has generally not led to improved functional predictions.
@@ I.1 A model experimental setup
The simplest, and most familiar, notion of emergent simplicity is self-averaging. As community richness increases, one might observe the distribution of $Y_{\mu}$ become increasingly tight. Much previous work suggests that certain properties of high-richness communities may indeed exhibit such “functional convergence.” [1, 2, 3, 4, 5, 6, 7, 8, 14] However, for the purposes of this work, what we care about is not the total variance of $Y_{\mu}$, but our ability to explain it from community composition. Our goal is to quantify the predictive power of simple models in ecological datasets, and then use empirical data to ask how this predictive power changes with community richness.
@@ I.2 The candidate ‘simple models’
Our approach is motivated by an empirical example, namely wastewater treatment bioreactors that use microbes to digest organic waste into methane. These complex communities assemble semi-naturally and can harbor over 1000 taxa [19]. The rate of digestion (the amount of methane produced) can vary widely [20, 21], but, remarkably, can be explained by surprisingly simple models coarse-graining this vast diversity into just 4 functional classes (acidogens, propionate-degrading acetogens, butyrate-degrading acetogens, and methanogens [22]).
Motivated by this example, throughout this work, our candidate simple models will be indexed by the choice of a coarsening $\Psi$, which maps individual strains into a smaller number of classes: $[1\dots S]\xrightarrow{\Psi}[1\dots K^{\Psi}]$ with $K^{\Psi}\leq S$ (Fig. 1A). Some natural choices of coarsening $\Psi$ could be to group strains based on taxonomy or certain functional traits, but we will allow any choice of $\Psi$ as a valid coarsening.
For any $\Psi$, we will assess the prediction error of the best model from a specified model class $\mathcal{M}$, taking as input the combined abundances $\tilde{n}^{\Psi}_{j\mu}=\sum_{\Psi(i)=j}n_{i\mu}$. This approach echoes other recent attempts to relate community-level properties to coarse-grained compositional features [23, 24]. The model class $\mathcal{M}$ could be, for instance, the class of Generalized Lotka-Volterra models, or some general-purpose ansatz (e.g., training a neural network). In this work, we will pick a particularly simple class of models, namely linear regressions $Y_{\mu}=\beta^{\Psi}_{0}+\sum_{i}\beta^{\Psi}_{i}\tilde{n}^{\Psi}_{i\mu}+\epsilon_{\mu}$ (where $\epsilon_{\mu}$ is the residual noise). This choice is motivated by two considerations. First, we will want to apply our framework to data, and as we will see, the limitations of existing datasets make it preferable to pick a model class trainable on fewest data points. Second, it was recently shown that simple regressions can in fact be surprisingly predictive of function [25], motivating us to proceed.
@@ I.3 Quantifying the performance of simple models
If we were to evaluate description complexity $I(\Psi)$ and the fraction of unexplained variance for all possible coarsenings $\Psi$, the result would be expected to look like Fig. 1C. Each point in the point cloud corresponds to one choice of $\Psi$, and describes the predictive power of one candidate simple model.
@@ II.1 Standard ecological models do not generically exhibit either form of emergent simplicity
Before we apply our framework to real data, it is helpful to establish whether observing the behaviors of Fig. 1D would be surprising. For this, we consider two classic models widely used in theoretical ecology: the Generalized Lotka-Volterra model (GLV) with a random interaction matrix, and a Consumer-Resource model (CRM) with a random resource consumption matrix.
@@
Figure 3: Emergent simplicity is observed in both available empirical datasets (Refs. [26] and [27]). A-E: The Pareto fronts in error-complexity plane, computed for five properties of interest. Lines show the fraction of variance left unexplained by the best model within a given tier of description complexity, when dataset is restricted to communities of low richness (gray), or high richness (black). At high richness, the Pareto front shifts down and to the left, indicating that simpler descriptions become more predictive (compare to Fig. 1C). Each point and error bar shows the median quartiles across 100 random 50-50 splits of the data into training and testing sets. G,H: Illustrating the result by plotting predicted versus measured butyrate concentration (the points on the Pareto front indicated by the arrowheads in A). At high richness, a simple compositional feature explains 85% of variance in butyrate production (panel G). In contrast, for low-richness communities, even the most predictive single feature explains only 49% of variance (panel F).
We make two observations. First, we see that in many cases, models with $I(\Psi)\gtrsim 3$ bits can capture $\gtrsim 75\%$ of variation in these simulated data. (Note that we simulated 1000 samples, and as a result there is no overfitting signature: the Pareto fronts show no uptick at high complexity). Even in the absence of simulated experimental noise, this level of performance was not guaranteed. Thus, Fig. 2 confirms that the model class of linear regressions, while simple, is sufficiently expressive to describe datasets structured by ecological interactions, at least those captured by these standard modeling frameworks.
Second, we see that in both cases, the Pareto fronts shift up and right, not down (Fig. 1E) or left (Fig. 1D). This analysis confirms our intuition that the emergent simplicity as defined here is not a behavior one would generically expect. Contrast this to Fig. 2C, where the same CRM simulation data are replotted using the absolute prediction error rather than relative. Here, the Pareto fronts exhibit a clear shift downwards. This is the phenomenon of ‘emergent reproducibility’ or functional convergence, where the distribution of $Y$ gets narrower as richness increases. The contrast with Fig. 2A, B highlights that both forms of emergent simplicity shown in Fig. 1D, E are distinct from this more familiar scenario, widely reported both empirically [1, 2, 3, 4, 5, 6, 7, 8] and in models [14, 24]. In Fig. 1D, E and Fig. 2A, B, the prediction error of $Y$ is always normalized to the overall width of the distribution of $Y$ across examples of the same richness tier. Our goal is to quantify forms of emergent simplicity going beyond emergent reproducibility, and our normalization explicitly removes this simpler effect.
@@ II.2 Empirical data exhibit emergent predictability
Furthermore, while we formulated our analysis in abstract terms such as the shape of a Pareto front, we can also confirm the result in more familiar terms by plotting the measured property against the prediction of the simplest models identified by our framework. Two examples are shown in Fig. 3F, G, which correspond to the points on the Pareto fronts labeled with arrowheads in Fig. 3A. At high richness, a simple two-parameter regression explains 85% of variance in butyrate production (Fig. 3G). In contrast, for low-richness communities, no model of comparable complexity can explain more than 49% of variance; the best-performing example is showin in Fig. 3F.
II.3 Emergent simplicity in functional space
So far we focused on asking how composition predicts functional properties. However, since Clark et al. measured four functional properties simultaneously, we can leverage this fact to also investigate whether these properties are predictive of each other.
In this dataset, the four functional measurements (the concentrations of butyrate, acetate, succinate and lactate produced by the community) define, for each community, a point in what we will call the “functional space”. Fig. 4A shows that at high community richness, these points are found along a lower-dimensional subspace. Indeed, for high-richness communities, 73% of variance is explained by PC1, compared to only 41% for low-richness communities (Fig. 4B). This dominant axis of variation is correlated with community pH (Fig. 4C), consistent with Ref. [26] identifying pH-mediated interactions as a key factor structuring these communities.
The emergent low dimensionality in the functional space (Fig. 4A) is distinct and independent from the emergent predictive power of simple compositional features (Fig. 3A-E). Thus, Fig. 4A provides a separate line of evidence indicating that, at least in these synthetic ecosystems, multiple aspects of community structure indeed become simpler and more predictable at high richness.
II.4 Mechanistic origin of emergent simplicity is unknown
@@
For specifically butyrate, the results of Clark et al. [26] provide some insight into the mechanism. Of the 25 strains in this study, five are capable of producing butyrate. However, in high-richness communities, butyrate production is dominated by one of them, Anaerostipes caccae (AC). And indeed, the first point on the high-richness Pareto front in Fig. 3A is the grouping {AC, all other strains}, and provides an excellent prediction for the amount of butyrate produced (Fig. 3G). In contrast, in low-richness mixtures, all five strains contribute to the production, and no single compositional feature is similarly predictive.
Precisely why a single producer becomes dominant is unclear. Clark et al. hypothesized that this transition is associated with AC shifting at high community richness to an alternative metabolic pathway, utilizing lactate rather than sugars, a physiological switch mediated by sensing pH and/or the depletion of key resources. Identifying candidate mechanistic explanations for the other properties, and validating these hypotheses, would require new experiments. It is conceivable that the observations of Fig. 3A-D are all a consequence of specific physiological features of individual community members. However, given that we see a common outcome across two independent datasets, it is intriguing to hypothesize that emergent predictability as quantified here might be more general than a mechanistic idiosyncrasy of specific taxa.
@@ III Discussion
In this work, we developed an information-theoretic framework to nuance and quantify the notion of “emergent simplicity” in ecosystems of increasing community richness (treated as the control parameter). Our framework enabled us to define and distinguish emergent predictability and emergent coarse-grainability, and to contrast both of these phenomena with the emergent reproducibility considered in much previous work. We then applied our framework to data from two separate experiments on synthetic microbial communities, and demonstrated that 4 out of 5 measured properties exhibit robust evidence of emergent predictability, whereby simple compositional descriptions become more predictive at higher richness.
We should note that in general, community richness (number of species) is a poor metric of community diversity, as it is overly sensitive to detection thresholds, and fails to account for the relatedness of community members [28]. This was not an issue for our work, as the synthetic communities considered here were assembled from a small pool of well-delimited taxa. For natural communities, this simple metric of diversity would need to be replaced by one of the better-suited metrics developed in the ecological literature [29, 30, 31].
@@
Our framework suggests a shift of viewpoint: instead of talking about simple and complex ecosystems (as quantified, e.g., by community richness), we should be distinguishing simple and complex properties of ecosystems. In the language we presented, this distinction is encoded in the shape of the Pareto front, as illustrated in Fig. 5. For properties of interest behaving as in Fig. 3, a comparatively coarse description may provide a “good-enough” prediction power at a fraction of the entropy of the fully microscopic description. We can call such properties coarse-grainable (Fig. 5, green). Conversely, for other properties one might find that useful prediction requires substantial microscopic detail (non-coarse-grainable; Fig. 5, red). Remarkably, in the datasets considered here, all measured properties proved coarse-grainable (compare Fig. 5 and Fig. 3A-E). Further, our key result can be summarized as saying that even as community complexity is increased, the prediction complexity of a given property need not do the same, and can even become simpler, as we quantitatively demonstrate.

### v2

@@ Abstract
A long-standing hope of theoretical ecology is that, despite this complexity, some patterns might nevertheless be predictable [1, 2, 3]. The search for emergent regularities in complex ecosystems is as old as ecology itself. For example, classical work uncovered reproducible trophic structures across taxonomically divergent communities [4] or a consistent ratio of predator species to prey species [5, 6]. More recently, moving beyond the patterns persisting despite high ecosystem diversity, it was proposed that some regularities could perhaps arise because of it [7, 8, 9], coining the term “emergent simplicity” [7, 10, 11, 12, 13]. This notion is at the foundation of the interface between statistical physics and the so-called large-$N$ ecology, explaining species abundance distributions and other patterns as some parameter becomes large [14, 15, 16, 17, 1, 8, 18, 19].
@@
Figure 1: A quantitative framework for interrogating emergent simplicity in data. A: A general experimental protocol with community richness (number of strains) as a control parameter. A fixed library of strains is subsampled to assemble communities of low richness (top) or high richness (bottom). Community composition (strain abundances, $n_{i}$) and some property of interest $Y$ are measured in each community. B: Our approach will be to train models predicting $Y$ based on combined abundances of groups of taxa. Each choice of a coarsening $\Psi$ defines a regression model (coefficients $\{\beta_{i}^{\Psi}\}$), whose performance we will evaluate. C: For each coarsening $\Psi$, we can evaluate the description complexity $I(\Psi)$ (the extent of coarsening; see text) and the out-of-sample prediction error, quantified as the fraction of variance unexplained (FVU) by the corresponding regression model. The cartoon illustrates the expected general shape of the scatter plot in these axes, evaluated for all possible coarsenings $\Psi$ (points). Our focus is on the shape of the Pareto front (thick line), characterizing the best predictive power achievable by models of a given complexity. As community richness increases, the phenomenon of emergent simplicity is identified by the behavior of the Pareto fronts, and can take one of two distinct forms: D: emergent coarse-grainability (horizontal arrow), or E: emergent predictability (vertical arrow). We stress that our normalization—focusing on the fraction of variance unexplained, equivalent to normalizing all observables to unit variance—ensures that both forms of emergent simplicity defined here are distinct from the simplest law-of-large-numbers intuition (a.k.a. central limit theorem or “self-averaging”); see text.
@@
The idea of emergent simplicity is very appealing, suggesting that functional predictions could perhaps be made directly in the complex natural context [20]. However, so far this hope remains out of reach. Most reported instances of “emergent simplicity” fall under the umbrella of functional convergence or self-averaging, focusing on reproducibility rather than prediction [4, 5, 6, 21, 7, 10, 22, 23, 11, 12, 13]. Unfortunately, such reproducibility is often approximate or limited to simplified conditions [24], and even when it holds, offers limited predictive power. As an illustration, consider the statement: “if the soil ecosystem is diverse enough, the rate of process $P$ will be $\approx\!\!X$ per gram of biomass”. This blanket expectation would likely be somewhat naïve, but even if it were true, it would leave key questions unanswered – such as how the process rate would be affected by a perturbation, or what intervention would help increase it. As a result, observing emergent regularities of this kind has generally not led to improved functional predictions.
@@ I.1 A model experimental setup
The simplest, and most familiar, notion of emergent simplicity is self-averaging. As community richness increases, one might observe the distribution of $Y_{\mu}$ become increasingly tight. Much previous work suggests that certain properties of high-richness communities may indeed exhibit such “functional convergence.” [4, 5, 6, 21, 7, 10, 22, 23, 11, 12, 13, 8] However, for the purposes of this work, what we care about is not the total variance of $Y_{\mu}$, but our ability to explain it from community composition. Our goal is to quantify the predictive power of simple models in ecological datasets, and then use empirical data to ask how this predictive power changes with community richness.
@@ I.2 The candidate ‘simple models’
Our approach is motivated by an empirical example, namely wastewater treatment bioreactors that use microbes to digest organic waste into methane. These complex communities assemble semi-naturally and can harbor over 1000 taxa [25]. The rate of digestion (the amount of methane produced) can vary widely [26, 27], but, remarkably, can be explained by simple models coarse-graining this vast diversity into just 4 functional classes (acidogens, propionate-degrading acetogens, butyrate-degrading acetogens, and methanogens [28]).
Motivated by this example, throughout this work, our candidate simple models will be indexed by the choice of a coarsening $\Psi$, which maps individual strains into a smaller number of classes: $[1\dots S]\xrightarrow{\Psi}[1\dots K^{\Psi}]$ with $K^{\Psi}\leq S$ (Fig. 1A). Some natural choices of coarsening $\Psi$ might be to group strains based on taxonomy or key traits, but we will allow any choice of $\Psi$ as valid.
For any coarsening $\Psi$, we will assess the prediction error of the best model from a specified model class $\mathcal{M}$, taking as input the combined abundances $\tilde{n}^{\Psi}_{j\mu}=\sum_{\Psi(i)=j}n_{i\mu}$. This approach echoes other recent attempts to relate community-level properties to coarse-grained compositional features [29, 30]. The model class $\mathcal{M}$ could be, for instance, the class of Generalized Lotka-Volterra models, or some general-purpose ansatz (e.g., training a neural network). In this work, we will pick a particularly simple class of models, namely linear regressions $Y_{\mu}=\beta^{\Psi}_{0}+\sum_{i}\beta^{\Psi}_{i}\tilde{n}^{\Psi}_{i\mu}+\epsilon_{\mu}$ (where $\epsilon_{\mu}$ is the residual noise). This choice is motivated by two considerations. First, we will want to apply our framework to data, and as we will see, the limitations of existing datasets make it preferable to pick a model class trainable on fewest data points. Second, it was recently shown that simple regressions can in fact be surprisingly predictive of ecosystem function [31], motivating us to proceed.
@@ I.3 Quantifying the performance of simple models
Established results of statistical learning theory indicate that if we could evaluate and plot the fraction of unexplained variance against the description complexity $I(\Psi)$ for all possible coarsenings $\Psi$, the result would be expected to look like Fig. 1C. Each point in the point cloud corresponds to one choice of $\Psi$, and describes the predictive power of one candidate simple model.
@@ II.1 Standard ecological models do not generically exhibit either form of emergent simplicity
Before we apply our framework to real data, it is important to establish whether observing the behaviors of Fig. 1D would be surprising [32]. For this, we consider two classic models widely used in theoretical ecology: the Generalized Lotka-Volterra model (GLV) with a random interaction matrix, and a Consumer-Resource model (CRM) with a random resource consumption matrix.
@@
Figure 3: Both available empirical datasets (Refs. [33] and [34]) exhibit emergent predictability. A-E: The Pareto fronts in error-complexity plane, computed for five properties of interest. Lines show the fraction of variance left unexplained by the best model within a given tier of description complexity, when dataset is restricted to communities of low richness (orange), or high richness (purple). At high richness, the Pareto front shifts down, indicating that simpler descriptions become more predictive (emergent predictability, Fig. 1E). Each point and error bar shows the median quartiles across 100 random 50-50 splits of the data into training and testing sets. G,H: Illustrating the result by plotting predicted versus measured butyrate concentration (the points on the Pareto front indicated by the arrowheads in A). At high richness, a simple compositional feature explains 85% of variance in butyrate production (panel G). In contrast, for low-richness communities, even the most predictive model of comparable complexity explains only 49% of variance (panel F).
We make two observations. First, we see that models with $I(\Psi)\gtrsim 3$ bits are generally able to capture $\gtrsim 75\%$ of variation in these simulated data. (Note that we simulated 1000 samples, and as a result there is no overfitting signature: the Pareto fronts show no uptick at high complexity). Even in the absence of simulated experimental noise, this level of performance was not guaranteed. Thus, Fig. 2 confirms that the model class of linear regressions, while simple, can be sufficiently expressive to describe datasets structured by ecological interactions. The one notable exception is the high-diversity GLV, where linear regressions provide no explanatory power.
Second, we see that in both cases, the Pareto fronts shift up and right, not down (Fig. 1E) or left (Fig. 1D). This analysis confirms our intuition that the emergent simplicity as defined here is not a behavior one would generically expect. Contrast this to Fig. 2C, where the same CRM simulation data are replotted using the absolute prediction error rather than relative. Here, the Pareto fronts exhibit a clear shift downwards. This is the phenomenon of ‘emergent reproducibility’ or functional convergence, where the distribution of $Y$ gets narrower as richness increases. The contrast with Fig. 2A, B highlights that both forms of emergent simplicity shown in Fig. 1D, E are distinct from this more familiar scenario, widely reported both empirically [4, 5, 6, 21, 7, 10, 22, 11, 12, 13] and in models [8, 30]. In Fig. 1D, E and Fig. 2A, B, the prediction error of $Y$ is always normalized to the overall width of the distribution of $Y$ across examples of the same richness tier. Our goal is to quantify forms of emergent simplicity going beyond emergent reproducibility, and our normalization explicitly removes this simpler effect.
@@ II.2 Empirical data exhibit emergent predictability
Furthermore, while we formulated our analysis in abstract terms such as the shape of a Pareto front, we can also confirm the result in more familiar terms by plotting the measured property against the prediction of the simplest models identified by our framework. Two examples are shown in Fig. 3F, G, which correspond to the points on the Pareto fronts labeled with arrowheads in Fig. 3A. At high richness, a simple two-parameter regression explains 85% of variance in butyrate production (Fig. 3G). In contrast, for low-richness communities, no model of comparable complexity can explain more than 49% of variance; the best-performing example is shown in Fig. 3F.
II.3 Mechanistic origin of emergent predictability
@@
At least for the system of Clark et al., the emergent predictability appears to be related to the dominant structuring role of pH, which is one-dimensional. While the data is insufficient to demonstrate this directly (independent measurements of pH were conducted for only a subset of communities), we can obtain strong indirect evidence by leveraging the fact that four functional properties were measured simultaneously. These four measurements (the final concentrations of butyrate, acetate, succinate and lactate produced) define, for each community, a point in what we will call the “functional space”. Fig. 4A shows that at high community richness, these points are found along a lower-dimensional subspace. Indeed, for high-richness communities, 73% of variance is explained by PC1, compared to only 41% for low-richness communities (Fig. 4B). This dominant axis of variation is strongly correlated with community pH (Fig. 4C), consistent with Ref. [33] identifying pH-mediated interactions as a key factor structuring these communities. This observation explains the sudden drop in the Pareto fronts observed in Fig. 3A-D. Indeed, any combination of species whose abundance is correlated with pH will explain a substantial fraction of variance in any of the observables that have a loading on the dominant principal component of Fig. 4. As a result, even a little compositional information can be sufficient to improve predictive power, which is what the Pareto front indicates.
This argument reinforces the conclusion of Fig. 2 that the shape of the Pareto fronts observed in Fig. 3 is not due simply to the number of interacting species being large, but reflects the biology of the underlying communities. Crucially, the analysis of Fig. 3A-D used no knowledge of pH. Thus, our framework may provide a prior-knowledge-agnostic way for discovering the existence of strong ecosystem-structuring factors.
For specifically Fig. 3A (butyrate), the results of Clark et al. [33] provide some additional mechanistic insight. Of the 25 strains in this study, five are capable of producing butyrate. However, in high-richness communities, butyrate production is dominated by one of them, Anaerostipes caccae (AC). And indeed, the first point on the high-richness Pareto front in Fig. 3A is the grouping {AC, all other strains}, and provides an excellent prediction for the amount of butyrate produced (Fig. 3G). In contrast, in low-richness mixtures, all five strains contribute to the production, and no single compositional feature is similarly predictive.
Clark et al. hypothesized that this transition is associated with AC shifting at high community richness to an alternative metabolic pathway, utilizing lactate rather than sugars, a physiological switch mediated by sensing pH and/or the depletion of key resources. Identifying candidate mechanistic explanations for the other properties, and validating these hypotheses, would require new experiments. It is conceivable that the observations of Fig. 3A-D are all a consequence of specific physiological features of individual community members. However, the common outcome observed across two independent datasets suggests that emergent predictability as quantified here may be more general than a mechanistic idiosyncrasy of specific taxa.
@@ III Discussion
In this work, we developed an information-theoretic framework to nuance and quantify the notion of “emergent simplicity” in ecosystems of increasing community richness (treated as the control parameter). Our framework enabled us to define and distinguish emergent predictability and emergent coarse-grainability, and to contrast both of these phenomena with the emergent reproducibility considered in much previous work. We applied our framework to data from two separate experiments on synthetic microbial communities, and demonstrated that four out of five measured properties exhibit robust evidence of emergent predictability, whereby simple compositional descriptions become more predictive at higher richness.
@@
In three of these four cases, the phenomenon appears to be linked to the dominant structuring role of pH in these communities. This observation echoes other recent examples where coarse-grained predictions are enabled by general physiological [10], metabolic [3], or toxicity-mediated constraints [35]. Our results indicate that these structuring factors may have higher explanatory power at higher diversity, suggesting the existence of a subtler form of functional convergence, where high diversity leads not to better absolute reproducibility, but to a stronger collapse onto a lower-dimensional subspace governed by metabolic or physiological constraints. Understanding this behavior will require integrating ongoing theoretical investigations of large random ecosystems with new models capable of incorporating the structuring role of such shared constraints [30].
Perhaps the most exciting implication of our framework is that it suggests a paradigm shift: instead of talking about simple and complex ecosystems (as quantified, e.g., by community diversity), we should be distinguishing simple and complex properties of ecosystems. In the language we presented, this distinction is encoded in the shape of the Pareto front, as illustrated in Fig. 5. For some properties of interest, one might find that useful predictions require substantial microscopic detail (non-coarse-grainable; Fig. 5, red). However, all the properties in Fig. 3 exhibit the opposite behavior, whereby a comparatively coarse description may provide a “good-enough” predictive power at a fraction of the entropy of the fully microscopic description (coarse-grainable; Fig. 5, green). Our key result can be summarized as saying that even as community complexity increases, the prediction complexity of a given property need not do the same, and can even become simpler, as we quantitatively demonstrate.
@@
Our approach has limitations. Defining emergent simplicity requires specifying a class of models, and here, due to the limited size of the two available datasets, our choice (linear regression) was particularly simple. However, our theoretical framework is more general; with sufficient data, the natural generalization of the Pareto fronts in Figs. 1-3 would be to evaluate the predictive power of machine-learning models as a function of the dimension of the compressed (latent-space) representation of community composition. Understanding the extent to which emergent predictability depends on the choice of model class remains a question for future work. We should also note that in general, community richness (number of species) is a poor metric of community diversity, as it is overly sensitive to detection thresholds, and fails to account for the relatedness of community members [36]. This was not an issue for our work, as the synthetic communities considered here were assembled from a small pool of well-delimited taxa. However, for natural communities, this simple metric of diversity would need to be replaced by one of the better-suited metrics developed in the ecological literature [37, 38, 39].